HUBBLE SPACE TELESCOPE DAILY REPORT #2176 PERIOD COVERED: 0000Z (UTC) 07/29/98 - 0000Z (UTC) 07/30/98 Daily Status Report as of 211/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7827 (The Structure of Circumbinary Material in Multiple T-Tauri Systems) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class system UZ-Tauri to study the structure of its circumbinary material. Recent studies of T-Tauri populations in nearby star forming clouds have shown that the majority of stars exist in multiple {binary or higher} systems. The observations were completed as planned, and no problems were reported. 1.4 Completed Two Sets of WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star cluster NGC-1818 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed WF/PC-2 6713 (Geometric Measurement of Galaxy Distance) The WF/PC-2 was to be used to make observations of supernova remnant SN1978K in order to make a distance calibration for the galaxy in which it is located. The supernova is being used as a powerful new geometric method for the determination of a galaxy's distance, by obtaining high spatial resolution imaging polarization of the candidate supernova light echoes. As described in yesterday's report and HSTAR 6601, the guide star acquisition for this proposal failed and the take data flag stayed down. 1.6 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was to be used to make pure parallel spectrographic observations of sky regions when another instrument is prime. As described in yesterday's report and HSTAR 6601, the guide star acquisition for all iterations of this proposal failed and the take data flag stayed down. 1.7 Completed Three Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was to be used to conduct a non-scripted parallel POMS test. As described in yesterday's report and HSTAR 6601, the guide star acquisition for all iterations of this proposal failed and the take data flag stayed down. 1.8 Completed Six Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in yesterday's report and HSTAR 6601, the guide star acquisition for the first iteration of this proposal failed and the take data flag stayed down. Then, as described in HSTAR 6603 and section 2.1, the acquisition for the fourth iteration of this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.9 Completed Two Sets of STIS/CCD 7969 (STIS CCD G750L Fringe Flats at Intermediate MSM Positions) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field fringing in the G750L grating at a number of wavelength settings. WAVECALs were also taken at the same gain and wavelength setting. As described in yesterday's report and HSTAR 6601, the guide star acquisition for the first iteration of this proposal failed and the take data flag stayed down. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.10 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make snapshot survey observations of the ultraluminous infrared galaxy IR02364-47. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. As described in yesterday's report and HSTAR 6602, the guide star acquisition for this proposal failed and the take data flag stayed down. 1.11 Completed STIS/MA1 7439 (H-Lyman Alpha Emission from the Upper Atmosphere of Uranus) The Space Telescope Imaging Spectrograph (MA1) was to be used to make H-Lyman Alpha observations of the emissions from the upper atmosphere of the planet Uranus. An excess of H-Lyman Alpha emission from Uranus' sunlit hemisphere has been detected by the IUE satellite since 1982, and some excess was confirmed with the Voyager 2 UVS during the 1986 encounter with Uranus. As described in yesterday's report and HSTAR 6602, the guide star acquisition for this proposal failed and the take data flag stayed down. 1.12 Completed WF/PC-2 7431 (Snapshot Survey of Microlensed Source Stars) The WF/PC-2 was used to make snapshot survey images of microlensed source star 97-BLG-45. Gravitational microlensing surveys have revealed an excess of lensing events over theoretical predictions toward both the Large Magellanic Cloud and the Galactic Bulge. The observations were completed as planned, and no anomalies were reported. 1.13 Completed NIC1/2/3 7221 (Imaging of Quasar Absorber Systems) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of quasar LBQS1210+1. These data should detect emission from the foreground galaxy which causes the damped Lyman-alpha absorption line systems in the spectra of the quasars. The observations were completed as planned, and no problems were reported. 1.14 Completed Three Sets of STIS/CCD 7910 (STIS Extragalactic Parallel Survey) The Space Telescope Imaging Spectrograph (CCD) was used to make extragalactic slitless spectrographic observations as part of the parallel high galactic latitude survey to search for emission line galaxies. Data was taken in parallel, and are not subject to proprietary time. As described in HSTAR 6603 and section 2.1, the acquisition for the second iteration of this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.15 Completed NIC/2 7879 (Nuclear Rings in the IR: Hidden Super Star Clusters) The Near Infrared Camera (Camera Two) was used to make observations of the galaxy NGC-1512 in order to search in the infrared for star bursting nuclear rings. The rings contain large populations of super star clusters similar to those recently discovered in other types of starburst systems. The observations were completed as planned, and no anomalies were reported. 1.16 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local Group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.17 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.18 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.19 Completed Four Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.20 Completed NIC/1/2 7864 (The Host Galaxies of Radio-Quiet High- Redshift Quasars) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio-quite, high red shifted quasar MZZ9592 and its host galaxy. The underlying causes of the rapid cosmological evolution of the quasar population are a mystery around which are woven many different strands of contemporary cosmology. As described in HSTAR 6603 and section 2.1, the acquisition this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 The acquisition at 210/1820Z defaulted to fine lock on FGS-3 only when FGS-1 did not acquire lock. HSTAR 6603 was written. The proposals described in 1.8, 1.14, and 1.20 may have been affected. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 14 Successful : 13 Per HSTAR 6604, the FHST Roll update at using FHSTs 1 and 2 211/000758Z failed due to FHST-1. The following full mode updates were successful. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH